What is the side length of a rectangular prism with a volume of 144 meters cubed a width of 2 meters and a height of 6 meters?
Answer by jim_thompson5910(35256)

V = LWH

144 = L*2*6

144 = L*12

144/12 = L

12 = L

L = 12

The length is 12 meters